Casey Nelson is a divisional manager for Pigeon Company. His annual pay raises are largely determined by his division’s return on investment (ROI), which has been above 23% each of the last three years. Casey is considering a capital budgeting project that would require a $5,620,000 investment in equipment with a useful life of five years and no salvage value. Pigeon Company’s discount rate is 19%. The project would provide net operating income each year for five years as follows:

Sales $ 5,000,000
Variable expenses 2,240,000
Contribution margin 2,760,000
Fixed expenses:
Advertising, salaries, and other
fixed out-of-pocket costs
$ 860,000
Depreciation 1,124,000
Total fixed expenses 1,984,000
Net operating income $ 776,000

Click here to view Exhibit 13B-1 and Exhibit 13B-2, to determine the appropriate discount factor(s) using tables.

Required:

1. What is the project’s net present value?

2. What is the project’s internal rate of return to the nearest whole percent?

3. What is the project’s simple rate of return?

4-a. Would the company want Casey to pursue this investment opportunity?

4-b. Would Casey be inclined to pursue this investment opportunity?
Answer #1
Net operating income 776000
Add: Depreciation 1124000
Net cash flows 1900000
1
Now Year 1 Year 2 Year 3 Year 4 Year 5
Investment cost -5620000
Net cash flows 1900000 1900000 1900000 1900000 1900000
Total cash flows -5620000 1900000 1900000 1900000 1900000 1900000
PV factor @ 19% 1 0.840 0.706 0.593 0.499 0.419
Present value of cash flows -5620000 1596000 1341400 1126700 948100 796100
Net present value 188300
2
PV factor internal rate of return=5620000/1900000 = 2.958
The PV factor 2.958 for 5 years is closest to 21%
Internal rate of return = 21%
3
Simple rate of return = Net operating income/Investment cost
Simple rate of return = 776000/5620000= 13.8%
4a
Yes, the company would want Casey to pursue this investment as Net Present value is positive
4b
No, Casey would not be inclined to pursue this investment as as his ROI will decrease